OpenStudy (anikate):

A driver in a car traveling at a speed of 75 km/h sees a cat 125 m away on the road. How long will it take for the car to accelerate uniformly to a stop in exactly 122 m? Answer in units of s.

OpenStudy (anikate):

any clue??

OpenStudy (anonymous):

@pooja195 Help me

OpenStudy (anonymous):

@pooja195 im new to helping

OpenStudy (anikate):

do u know the answer??

OpenStudy (anikate):

@DylanisNOW u know the answer??

OpenStudy (anonymous):

Im trying to get @pooja195

OpenStudy (anikate):

i thuink she offline, but do know the answer to this question??

OpenStudy (anikate):

she appeears to be offline??

OpenStudy (anikate):

yea she offline

OpenStudy (anonymous):

Hold on...Im trying to figure it out. Please dont rush!

OpenStudy (anikate):

ok

OpenStudy (anonymous):

@pooja195 No she isnt

OpenStudy (anikate):

ok, do u know how to do it though?

OpenStudy (anonymous):

Give me a sec please

OpenStudy (anikate):

ok np

OpenStudy (anikate):

@pgpilot326 do u know how to do this? its a physics question

OpenStudy (anonymous):

Think of the units.

OpenStudy (anikate):

ok now?

OpenStudy (anonymous):

units are 75km/h right?

OpenStudy (anikate):

yes

OpenStudy (anonymous):

Wait sorry hold on thats not it thats only 1

OpenStudy (anikate):

ok

OpenStudy (anonymous):

75km/h / 125 m=?

OpenStudy (anikate):

600m/h

OpenStudy (anonymous):

Okay do you have answer choices?

OpenStudy (anikate):

nopee

OpenStudy (anikate):

do u havea roungh estimate of what the andswer could be??

OpenStudy (anonymous):

Think of it like this. Every second the speed decreases by?

OpenStudy (anonymous):

We cant give answers...Its just help to get you to understand it!

OpenStudy (anonymous):

Think of it like this. Every second the speed decreases by?

OpenStudy (anikate):

the speed decreases by seconds(squared)

OpenStudy (anonymous):

Ok that is?

OpenStudy (anikate):

idk.....

OpenStudy (anonymous):

.... @pgpilot326 can you help me help him?

OpenStudy (anonymous):

Anyways..

OpenStudy (anikate):

dylanis, can u just post a step by step explaination for it

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Step 1:units are 75km/h

OpenStudy (anonymous):

Step 2:75km/h / 125=600m/h

OpenStudy (anonymous):

Okay now we are at step 3

OpenStudy (anonymous):

if a is a constant, integrate to get velocity and again to get position. the velocity should be linear and the position a quadratic.

OpenStudy (anonymous):

since you know the initial and final speed you should be able to calc. the time

OpenStudy (anonymous):

pgpilot should we make a equation?

OpenStudy (anonymous):

should make 2 equations... velocity and position. you'll need both.

OpenStudy (anikate):

can you please show bothh?

OpenStudy (anikate):

and solve with steps please?

OpenStudy (anonymous):

\[v \left( t \right)=v_0+at\]and\[x \left( t \right)=x_0+v_0t+\frac{ 1 }{ 2 }at^2\]

OpenStudy (anonymous):

pg you are right good job btw :)

OpenStudy (anonymous):

at t=0, initial speed (\(v_0\)) is 75 km/h and initial position is 0 m. at t=\(t_1\), final speed (\(v_f\)) is 0 km/h and final position is 122 m.

OpenStudy (anonymous):

you need to solve for \(t_1\)

OpenStudy (anikate):

ok

OpenStudy (anikate):

how so??

OpenStudy (anikate):

which equation am i putting all this in??

OpenStudy (anonymous):

phone, just a minute

OpenStudy (anonymous):

\[v \left( t_1 \right)=0=75 +a\cdot t_1\] \[x \left( t_1 \right)=122=0+75t_1+\frac{ 1 }{ 2 }at_1^2\]

OpenStudy (anonymous):

find roots of quadratic in terms of a.

OpenStudy (anikate):

u can handwrite it if u wish, ill prob be able to decode it

OpenStudy (anikate):